OpenStudy (anonymous):

solve equation for x: 4x-8y=16

OpenStudy (theviper):

Where is the other equation ??

OpenStudy (anonymous):

good question where is the whole problem it make no sense otherwise

OpenStudy (anonymous):

4x-8y-16

OpenStudy (anonymous):

4x-8y=16 solve x

OpenStudy (anonymous):

So you need to get x alone so start moving terms away from x. \[4x-8y=16\] Add "8y" to each side to get rid of the 8y on the left: \[4x-8y+8y=16+8y\]\[4x=16+8y\] Divide each term by 4 to get rid of the 4 on the left: \[4x/4 = 16/4 + 8y/4\]\[x=4+2y\]

OpenStudy (anonymous):

ok thanks, i have the same answer